Mathematics 351 – Abstract Algebra 1 Solutions for Problem Set 9 November 9, 2007

Section 7.5

5. (a) Z_{24} has order 24, so by Lagrange’s Theorem, the possible orders of subgroups of Z_{24} are 1, 2, 3, 4, 6, 8, 12, 24.